OpenStudy (anonymous):

Help :) Solve the equation. Log(x+7)- Logx=3 A. 3.5 B. 142.7 C. 0.0070 D. 0.0707

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

properties of logs:\(\log_b(\dfrac{m}{n}) = \log_b(m) – \log_b(n)\)

OpenStudy (anonymous):

So would i do logb (7/3)= logb (7)- logb (3)?

OpenStudy (bibby):

here, b is 10 \(\log(x+7)- \log x=3\\\log(\dfrac{x+7}{x})=3\\10^3=\dfrac{x+7}{x}\\1000x=x+7\\999x=7\\x=\dfrac{7}{999} \)